OpenStudy (anonymous):

2sinx+(1-sinx)^2=2-cos^2x Prove identity Will give medal

OpenStudy (trojanpoem):

L.H.S = 2sinx + (1- sinx)^2 = 2sinx + 1 + sin^2x -2sinx = 1 + sin^2x -> (1) sin^2x + cos^2x = 1 sin^2x = 1 - cos^2x ->(2) from both (1), (2); L.H.S = 1 + 1 - cos^2x = 2- cos^2x = R.H.S #